[figure: 9.7 cm]

MAGNETOELASTIC ANOMALIES AND SUPPRESSION OF JAHN-TELLER DISTORTIONS IN La_{1−x}Sr_{x}MnO_{3} BY PULSED

MAGNETIC FIELDS

A.M. Kadomtseva ^{a}, Yu.F. Popov ^{a}, G.P. Vorobev ^{a}, K.I. Kamilov ^{a}, V.Yu. Ivanov ^{b}, A.A. Mukhin ^{b}, A.M. Balbashov ^{c} aMoscow State University, Leninskie Gori, 119899 Moscow, Russia bGeneral Physics Institute of the RAS, 38 Vavilov St., 117942 Moscow, Russia cMoscow Power Engineering Institute, 14 Krasnokazarmennaya St., 105835 Moscow, Russia

Thermal expansion Δl/l(T) and magnetostriction λ_{∥,⊥}(H, T) in the pulsed magnetic fields H up to 250 kOe were studied in the single La_{1−x}Sr_{x}MnO_{3} crystals (x=0.1, 0.125 and 0.15) at temperatures T =10-320 K. Anomalies in the Δl/l(T) curves were observed near a charge (polaron) ordering (T_{p}) and structural transitions between orthorhombic Jahn-Teller distorted O´ phase and high temperature orthorhombic (pseudocubic) O^{∗} one (T_{s}´ ). Jumps in magnetic field dependence of the magnetostriction were observed in a narrow temperature range between charge and magnetic orderings indicating on an existence of field induced transitions, which threshold fields were strongly increased with the temperature. Noticeable hysteric phenomena accompanied the magnetostriction behavior. A noticeable increase of the field-induced magnetostriction λ_{∥,⊥}(H) (from 10^{−4} up to 7·10^{−4}) was observed in the O´ phase at T_{p} < T < T_{s}´ for the all compositions studied. The observed features may be related to a suppression of the Jan-Teller O´ phase by the magnetic field and field induced transition to an orbital-ordered ferromagnetic state. This work was supported in part by the RFBR (N 97-02-17325).

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : Vsevolod Yu. Ivanov

Address for correspondence : General Physics Institute of the Russian Acad. Sci., Vavilov St., 38. 117942 Moscow, Russia

Email address : ivanov@ran.gpi.ru